# Learning Partial Differential Equations with Deep Parallel Neural Operator

### v2

@@ 4.2.6 Pipe problem with Navier-Stokes equation
where $\mathbf{v}$ is the velocity vector, $p$ is the pressure, and $\nu=0.005$ is the viscosity. In this case, at the inlet, a parabolic velocity profile is established with a maximum velocity of $\mathbf{v}=[1;0]$. The outlet is treated with a free boundary condition, while the pipe surface is subject to a no-slip boundary condition. The pipe has a length of 10 and a diameter of 1. The centerline of the pipe is defined using four piecewise cubic polynomials, which are derived from the vertical positions and slopes at five evenly spaced control nodes. The vertical positions at these nodes follow a uniform distribution $d\sim U[-2;2]$, and the slopes at these nodes are also distributed uniformly, $d\sim U[-1;1]$ [28]. The experiment uses 1000 training grid points and 200 test grid points. The results are shown in Table 9 and Fig 6.
Table 9: Comparison of Models for Pipe Problem
@@
DPNO 0.0052(-3.8%)
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 6: Pipe Problem.
@@ 4.3 Results and Discussion
DPNO achieved state-of-the-art performance on five benchmark datasets and one sub-optimal. An average improvement of 10.5% is achieved compared to the overall best performing LSM. This demonstrates the versatility and excellent performance of the proposed DPNO in different PDE tasks. Compared to the baseline model FNO, DPNO shows an average performance improvement of 39% on the six datasets. Among them, the performance of the PLAYCITY-3d problem is improved by 74%. As shown in table 10:
Table 10: Comparison of MSE for the 7 benchmark models on the 6 benchmark datasets. The best performing models are in bold, and the next best models are underlined.
@@ 4.4 Ablation Experiment
(a) Parallel Operator Block.
(b) Serial Operator Block.
Figure 7: Comparison of Operator Block Architectures.
The experimental results are shown in Table 11:
@@
DPNO 0.0052 0.0054 0.0375 0.0019 0.0062 0.1418 9.3%
Through the ablation experiments, we can find that parallelisation has some degree of impact on the overall model performance, with an average performance improvement of 9.3%. This is because by parallelising the module, we mitigate the error accumulation caused by the iterations.
@@ 5 Conclusions
Neural operators are good learners when it comes to functional spaces. In this paper, we propose a deeply parallel Fourier neural operator to address the problem of error accumulation during FNO iterations. The data are projected into various potential spaces and learnt in an end-to-end manner through multiple parallel operator blocks. Excellent performance is achieved on six benchmark datasets for partial differential equations. An average improvement of 39% is achieved compared to the baseline model FNO.
@@ Appendix A Navier-stokes Experiment
Where the first row is the model prediction, the second row is the absolute value error. The Navier-stokes equation used in this paper is a task that includes a time term. Each column in the image represents a state at a different point in time.
(a) Experimental results of DPNO on Navier-stokes Equation.
(b) Experimental results of FNO on Navier-stokes Equation
(c) Experimental results of LSM on Navier-stokes Equation.

### v3

@@ 4.2.6 Pipe problem with Navier-Stokes equation
where $\mathbf{v}$ is the velocity vector, $p$ is the pressure, and $\nu=0.005$ is the viscosity. In this case, at the inlet, a parabolic velocity profile is established with a maximum velocity of $\mathbf{v}=[1;0]$. The outlet is treated with a free boundary condition, while the pipe surface is subject to a no-slip boundary condition. The pipe has a length of 10 and a diameter of 1. The centerline of the pipe is defined using four piecewise cubic polynomials, which are derived from the vertical positions and slopes at five evenly spaced control nodes. The vertical positions at these nodes follow a uniform distribution $d\sim U[-2;2]$, and the slopes at these nodes are also distributed uniformly, $d\sim U[-1;1]$ [28]. The experiment uses 1000 training mesh points and 200 test mesh points. For this problem, neural networks are mainly concerned with how to model the velocity field of a fluid through geometric parameters. Grid point location parameters (129*129) and horizontal velocity on these mesh points are used as input and output data. The results are shown in Table 9 and Fig 6.
Table 9: Model performance comparison on the Pipe Problem benchmark.
@@
DPNO 0.0052
Figure 6: Pipe Problem: The first row shows the true values and model predictions, the second row shows the model errors, and the legends are the true value legend and the error legend, respectively.
@@ 4.3 Results and Discussion
DPNO achieved state-of-the-art performance on five benchmark datasets and one sub-optimal. An average improvement of 11.6% is achieved compared to the overall best performing LSM. This demonstrates the versatility and excellent performance of the proposed DPNO in different PDE tasks. Compared to the baseline model FNO, DPNO shows an average performance improvement of 40% on the six datasets. Among them, the performance of the PLAYCITY-3d problem is improved by 73%. As shown in table 10: The comparison chart of model training errors stability can be found in B.
Table 10: Comparison of MSE for the 7 benchmark models on the 6 benchmark datasets. The best performing models are in bold, and the next best models are underlined. Promotion refers to the relative error reduction w.r.t. the second best model on each benchmark.
@@ 4.4 Ablation Experiment
(a) Parallel Fourier Block.
(b) Serial Fourier Block.
Figure 7: Comparison of Fourier Block Architectures.
Due to the unidirectional nature of the serial, taking too small a modal truncation factor can lead to too much lost information, therefore, Serial Fourier blocks use the same modal truncation coefficients for both Fourier layers. Low frequency information tends to carry more physical information, The parallel Fourier block utilises a decreasing strategy for the modal truncation coefficients, this is due to the consistency of the inputs to each Fourier layer, changes in the modal truncation coefficients do not affect the respective learning process. The experimental results are shown in Table 11:
@@
DPNO 0.0052 0.0054 0.0375 0.0020 0.0057 0.1303 10.8%
Through the ablation experiments, we can find that parallelisation has some degree of impact on the overall model performance, with an average performance improvement of 10.8%. This is because by parallelising the module, we mitigate the error accumulation caused by the iterations.
@@ 5 Conclusions
Neural operators are good learners when it comes to functional spaces. In this paper, we propose a deeply parallel Fourier neural operator to address the problem of error accumulation during FNO iterations. The data are projected into various potential spaces and learnt in an end-to-end manner through multiple parallel operator blocks. Excellent performance is achieved on six benchmark datasets for partial differential equations. An average improvement of 40% is achieved compared to the baseline model FNO.
@@ Appendix A Navier-stokes Experiment
Where the first row represents the true value, the second row is the model predicted value and the third row is the error. Each column corresponds to a point in time.
@@
Appendix B Training Stability
We provide the training error curves of the model on different benchmarks in Fig 9. From Fig 9, we can see that in addition to maintaining state-of-the-art performance on all benchmarks, there is comparable training stability and convergence speed compared to the SOTA model LSM and the recognised FNO.
In addition, for one experiment, we repeated the experiment at least five times or more, and all errors were within 0.0005.
Figure 9: Model training curves, where the x-axis means the number of epochs and the y-axis is MSE performance in the test set. The inset on the right shows a zoomed-in view of the training after 400 epochs.